# Couples can be tractable: New algorithms and hardness results for the Hospitals / Residents problem with Couples

### v2

@@ 4.2 Dual markets
Suppose not. Then $y_{4i+r}$ is either unmatched in $M^{\prime}$ or $M^{\prime}(y_{4i+r})=u^{5}_{4i+r}$. If $y_{4i+r}$ is unmatched then $(u^{5}_{4i+r},y_{4i+r})$ blocks $M^{\prime}$ in $I^{\prime}$, a contradiction. Hence, $M^{\prime}(y_{4i+r})=u^{5}_{4i+r}$ and thus $(u^{5}_{4i+r},h^{1}_{4i+r})\notin M$.
@@
We now show through the following two lemmata that if $M^{\prime}$ is a stable matching in $I^{\prime}$ then $M=M^{\prime}\setminus(U\times H)$ is a complete stable matching in $I$.
@@
Assume $z^{1}_{j}$ is unmatched in $M^{\prime}$ for some $j$ $(1\leq j\leq m)$. Thus $(p_{j}^{s},z^{2}_{j})\notin M$ $(4\leq s\leq 6)$ as $z^{2}_{j}$ must also be unmatched. Hence, $(z^{1}_{j},z^{2}_{j})$ are jointly unmatched and find $(p^{1}_{j},p^{4}_{j})$ acceptable. Further, $(p^{1}_{j},p^{4}_{j})$ jointly prefers $(z^{1}_{j},z^{2}_{j})$ to any other pair. Hence $(z^{1}_{j},z^{2}_{j})$ blocks $M$ with $(p^{1}_{j},p^{4}_{j})$, a contradiction. Thus, $z^{1}_{j}$ must be matched in any stable matching admitted by $I^{\prime}$. By a similar argument, the same holds for $z^{2}_{j}$.
@@
Let $B$ be an instance of (2,2)-e3-sat. Construct an instance $I$ of hrc-dual market as described in the proof of Theorem 22 and extend this instance as described above by adding the enforcer gadgets corresponding to the doctors in $U$ and the hospitals in $H$, to obtain the instance $I^{\prime}$ of hrc.
@@
As shown in the proof of Theorem 22, no agent outside of $Y\cup U\cup H$ can block $M^{\prime}$ in $I$. For each $i$ ($10\leq i\leq n-$ 1) and $r$ ($0\leq r\leq 3$), since $h_{4i+r}^{1}$ has its first-choice doctor in $M^{\prime}$, the couple $(u^{1}_{4i+r},u^{2}_{4i+r})$ cannot block $M^{\prime}$ in $I^{\prime}$. For a similar reason, the couple $(u^{3}_{4i+r},u^{4}_{4i+r})$ cannot block $M^{\prime}$ in $I$. Finally $M^{\prime}(y_{4i+r})\in X$, hence $y_{4i+r}$ cannot block $M^{\prime}$ in $I^{\prime}$. Thus $M^{\prime}$ is stable in $I^{\prime}$.
@@
Finally, the master lists shown in Figures 11, 12 and 13 indicate that the preference list of each single doctor, couple and hospital may be derived from a master list of hospital pairs, doctors and hospitals, respectively. ∎
$\begin{array}[]{c}L^{1}_{i}:(y_{4i},l_{4i}),(c(x_{4i}),l_{4i+1}),(y_{4i+1},l_{4i+1}),(c(x_{4i+1}),l_{4i+2}),(c(x_{4i+3}),l_{4i+3}),(y_{4i+3},l_{4i+3}),(c(x_{4i+2}),l_{4i+2}),(y_{4i+2},l_{4i+2})~{}~{}~{}~{}(0\leq i\leq n-1)\\ \\ L^{2}_{i,r}:(h^{1}_{4i+r},h^{2}_{4i+r}),(h^{1}_{4i+r},h^{4}_{4i+r}),(h^{3}_{4i+r},h^{2}_{4i+r})~{}~{}~{}~{}(0\leq i\leq n-1,0\leq r\leq 3)\par~{}~{}~{}~{}~{}~{}~{}~{}~{}\par L^{3}_{j}:(z^{1}_{j},z^{2}_{j}),(c^{1}_{j},z^{3}_{j}),(c^{2}_{j},z^{4}_{j}),(c^{3}_{j},z^{5}_{j})~{}~{}~{}~{}(1\leq j\leq m)\\ \\ \mbox{Master List}:L^{1}_{0}~{}L^{1}_{1}~{}\dots~{}L^{1}_{n-1}~{}L^{2}_{0,0}~{}L^{2}_{0,1}~{}L^{2}_{0,2}~{}L^{2}_{0,3}~{}L^{2}_{1,0}~{}L^{2}_{1,1}~{}L^{2}_{1,2}~{}L^{2}_{1,3}\ldots L^{2}_{(n-1),0}~{}L^{2}_{(n-1),1}~{}L^{2}_{(n-1),2}~{}L^{2}_{(n-1),3}~{}L^{3}_{1}~{}L^{3}_{2}~{}\dots~{}L^{3}_{m}\par\end{array}$
@@
$\begin{array}[]{c}L^{4}_{j}:c^{1}_{j},c^{2}_{j},c^{3}_{j},z^{3}_{j},z^{4}_{j},z^{5}_{j}\par~{}~{}~{}~{}(1\leq j\leq m)\par~{}~{}~{}~{}~{}~{}~{}~{}\par L^{5}_{i,r}:\par y_{4i+r},h^{1}_{4i+r}~{}~{}~{}~{}(0\leq i\leq n-1,0\leq r\leq 3)\\ \\ \mbox{Master List}:L^{4}_{1}~{}L^{4}_{2}~{}\dots~{}L^{4}_{m}~{}L^{5}_{0,0}~{}L^{5}_{0,1}~{}L^{5}_{0,2}~{}L^{5}_{0,3}~{}L^{5}_{1,0}~{}L^{5}_{1,1}~{}L^{5}_{1,2}~{}L^{5}_{1,3}~{}\ldots L^{5}_{(n-1),0}~{}L^{5}_{(n-1),1}~{}L^{5}_{(n-1),2}~{}L^{5}_{(n-1),3}\end{array}$
@@
$\begin{array}[]{c}L^{6}_{i}:x_{4i+1},x_{4i},x_{4i+2},x_{4i+3},k_{4i+3},k_{4i},k_{4i+2},k_{4i+1}~{}~{}~{}~{}(0\leq i\leq n-1)\\ \\ L^{7}_{j}:p^{1}_{j},p^{2}_{j},p^{3}_{j},p^{6}_{j},p^{5}_{j},p^{4}_{j}~{}~{}~{}~{}(1\leq j\leq m)~{}~{}~{}~{}~{}~{}~{}~{}L^{8}:s_{1},s_{2},\dots,s_{m},t_{1},t_{2},\dots,t_{m}\\ \\ L^{9}_{i,r}:u^{5}_{4i+r},u^{1}_{4i+r},u^{3}_{4i+r},u^{4}_{4i+r},u^{2}_{4i+r}~{}~{}~{}~{}(0\leq i\leq n-1,0\leq r\leq 3)\\ \\ \mbox{Master List}:L^{7}_{1}~{}L^{7}_{2}~{}\dots~{}L^{7}_{m}~{}L^{6}_{0}~{}L^{6}_{1}~{}\dots~{}L^{6}_{n-1}~{}L^{8}~{}L^{9}_{0,0}~{}L^{9}_{0,1}~{}L^{9}_{0,2}~{}L^{9}_{0,3}~{}L^{9}_{1,0}~{}L^{9}_{1,1}~{}L^{9}_{1,2}~{}L^{9}_{1,3}~{}\ldots~{}L^{9}_{(n-1),0}~{}L^{9}_{(n-1),1}~{}L^{9}_{(n-1),2}~{}L^{9}_{(n-1),3}\end{array}$
@@ 4.3 Matchings with the minimum number of blocking pairs
min bp hrc is not approximable within $m^{1-\varepsilon}$, for any $\varepsilon>0$, where $m$ is the total length of the hospitals’ preference lists, unless P=NP. The result holds even if there are no single doctors, every couple finds acceptable only one pair of hospitals, and each hospital has capacity 1.
@@
Let $\varepsilon>0$ be given. We reduce from com smti, which is shown to be NP-complete in the Appendix of [25] even if every man’s list is strictly ordered and of length at most 3, and every woman’s list is either strictly ordered and of length at most 3, or is a tie of length 2. Hence let $I$ be an instance of this restriction of com smti.
Let $U$ be the set of men in $I$ and let $W$ be the set of women in $I$. Let $G=(U\cup W,E)$ denote the underlying bipartite graph of $I$, and let $m_{I}$ denote $|E|$. Let $W^{s}$ denote the set of women in $W$ whose preference lists are strictly ordered in $I$, and let $W^{t}=W\setminus W^{s}$. Then every woman in $W^{t}$ has a preference list in $I$ that is a tie of length 2. Let $n_{t}=|W^{t}|$. Let $C=\lceil\frac{3}{\varepsilon}\rceil$ and let $B=n_{t}m_{I}^{C}+1$.
We construct an instance $J$ of hrc on the basis of $I$ as follows. Let $H$ be the set of hospitals in $J$, where $H=H_{U}\cup H_{W}$, $H_{U}=\{h_{m}:m\in U\}$ and $H_{W}=\{h_{w}:w\in W\}$. Let $H_{W}^{s}=\{h_{w}:w\in W^{s}\}$, and let $H_{W}^{t}=H\setminus H_{W}^{s}$.
For each person $p\in U\cup W$, let $\succeq_{p}$ denote $p$’s preference list in $I$. Each hospital $h_{p}\in H$ is initially given the preference list $\succeq_{p}$ in $J$. Note that if $p\in W^{t}$ then the tie of length 2 is broken arbitrarily in $\succeq_{p}$ so that we obtain a strictly ordered ranking $\succ_{p}$ for every agent $h_{p}$ in $J$.
Now let $e=\{m,w\}\in E$. Suppose firstly that $w\in W^{t}$. Create the couple $C^{e}=(c_{m}^{e},c_{w}^{e})$ who find acceptable the hospital pair $(h_{m},h_{w})$. Replace $m$ by $c_{m}^{e}$ in $\succ_{h_{w}}$, and replace $w$ by $c_{w}^{e}$ in $\succ_{h_{m}}$. Now suppose that $w\in W^{s}$. Create the couple $C^{e,k}=(c_{m}^{e,k},c_{w}^{e,k})$ who find acceptable the hospital pair $(h_{m},h_{w})$, for each $k$ ($1\leq k\leq B$). Replace $m$ by $c_{m}^{e,1}\succ_{h_{w}}\dots\succ_{h_{w}}c_{m}^{e,B}$ in $\succ_{h_{w}}$, and replace $w$ by $c_{w}^{e,1}\succ_{h_{m}}\dots\succ_{h_{m}}c_{w}^{e,B}$ in $\succ_{h_{m}}$. Let $P(h_{w})$ denote $\succ_{h_{w}}$ at this point.
@@
Every hospital in $J$ has capacity 1. It is easy to see that there are no single doctors in $J$, and each couple finds acceptable only one pair of hospitals. We claim that (i) if $I$ admits a complete stable matching then $J$ admits a matching with at most $n_{t}$ blocking pairs, and (ii) if $I$ admits no complete stable matching then any matching in $J$ has at least $B$ blocking pairs.
@@
To show (i), suppose that $M$ is a complete stable matching in $I$. Construct a matching $M^{\prime}$ in $J$ as follows. For each pair $(m,w)$ in $M^{\prime}$, representing the underlying edge $e$ in $G$, suppose firstly that $w\in W^{t}$. Match the couple $(c_{m}^{e},c_{w}^{e})$ with the hospital pair $(h_{m},h_{w})$ in $M^{\prime}$. Now suppose that $w\in W^{s}$. Match the couple $(c_{m}^{e,1},c_{w}^{e,1})$ with the hospital pair $(h_{m},h_{w})$ in $M^{\prime}$. Finally, for each $h_{w}\in H_{W}$ and $k$ ($1\leq k\leq B$), match the couple $(c_{w,3}^{k},c_{w,4}^{k})$ with the hospital pair $(h_{w}^{\prime},h_{w}^{\prime\prime})$. Each hospital is then full in $M^{\prime}$, since $M$ is a complete stable matching in $I$.
@@
For each $h_{w}\in H_{W}$, and $k$ ($1\leq k\leq B$), it may be verified that $(c_{w,1}^{k},c_{w,2}^{k})$ cannot be involved in blocking $M^{\prime}$ in $J$ , since $h_{w}$ is assigned to a better doctor than $c_{i,1}^{1}$ in $M^{\prime}$. Similarly $(c_{w,5}^{k},c_{w,6}^{k})$ cannot be involved in blocking $M^{\prime}$ in $J$ for the same reason. Indeed, since $M$ is a complete stable matching in $I$, the only possible blocking pairs of $M^{\prime}$ in $J$ involve couples of the form $(c_{m}^{e},c_{w}^{e})$, where $w\in W^{t}$. Since each woman in $W^{t}$ has a preference list of length 2 in $J$, there can be at most one blocking pair for each $w\in W^{t}$, and hence $|bp(J,M^{\prime})|\leq n_{t}$.
@@
To show (ii), suppose that $I$ admits no complete stable matching, and let $M^{\prime}$ be any matching in $J$. Suppose firstly that some hospital $h_{w}$ is unassigned in $M^{\prime}$ ($h_{w}\in H_{W}$). Then $(c_{w,5}^{k},c_{w,6}^{k})$ is unassigned in $M^{\prime}$ for all $k$ ($1\leq k\leq B$). Hence $(c_{w,5}^{k},c_{w,6}^{k})$ blocks $M^{\prime}$ with $(h_{w}^{\prime\prime},h_{w})$ for all $k$ ($1\leq k\leq B$), so $|bp(J,M^{\prime})|\geq B$. Now suppose that $(c_{w,6}^{k},h_{w})\in M^{\prime}$ for some $h_{w}\in H_{W}$ and $k$ ($1\leq k\leq B)$. Then $(c_{w,5}^{k},h_{w}^{\prime\prime})\in M^{\prime}$. Hence $(c_{w,1}^{k^{\prime}},c_{w,2}^{k^{\prime}})$ and $(c_{w,3}^{k^{\prime}},c_{w,4}^{k^{\prime}})$ are unassigned in $M^{\prime}$ for all $k^{\prime}$ ($1\leq k^{\prime}\leq B)$ and also $h_{w}^{\prime}$ is unassigned in $M^{\prime}$. Hence $(c_{w,1}^{k^{\prime}},c_{w,2}^{k^{\prime}})$ blocks $M^{\prime}$ with $(h_{w},h_{w}^{\prime})$ for all $k^{\prime}$ ($1\leq k^{\prime}\leq B$), so $|bp(J,M^{\prime})|\geq B$. Finally suppose that $(c_{w,1}^{k},h_{w})\in M^{\prime}$ for some $h_{w}\in H_{W}$ and $k$ ($1\leq k\leq B)$. Then $(c_{w,2}^{k},h_{w}^{\prime})\in M^{\prime}$. Hence $(c_{w,3}^{k^{\prime}},c_{w,4}^{k^{\prime}})$ and $(c_{w,5}^{k^{\prime}},c_{w,6}^{k^{\prime}})$ are unassigned in $M^{\prime}$ for all $k^{\prime}$ ($1\leq k^{\prime}\leq B)$ and also $h_{w}^{\prime\prime}$ is unassigned in $M^{\prime}$. Hence $(c_{w,3}^{k^{\prime}},c_{w,4}^{k^{\prime}})$ blocks $M^{\prime}$ with $(h_{w}^{\prime},h_{w}^{\prime\prime})$ for all $k^{\prime}$ ($1\leq k^{\prime}\leq B$), so $|bp(J,M^{\prime})|\geq B$.
@@
We thus assume that each $h_{w}\in H_{W}$ has an assignee better than $c_{w,1}^{1}$ in $M^{\prime}$. Let $M$ be the matching in $I$ obtained by replacing each assignment of a couple $(c_{m}^{e},h_{w}^{e})$ with the pair $(m,w)$ in $M$, and by replacing each assignment of a couple $(c_{m}^{e,k},h_{w}^{e,k})$ with the pair $(m,w)$ in $M$. It is easy to see that $M$ is a matching in $I$, since each of $(c_{m}^{e},h_{w}^{e})$ and $(c_{m}^{e,k},h_{w}^{e,k})$ finds acceptable only $(h_{m},h_{w})$, and each hospital has capacity 1 in $J$. Moreover $M$ is a complete matching in $I$, since in $M^{\prime}$, each hospital $h_{w}\in H_{W}$ has an assignee better than $c_{w,1}^{1}$. Hence by assumption, $M$ is not stable in $I$, so $M$ admits a blocking pair $(m,w)$ in $I$.
@@
Let $e=(m,w)\in E$. Since the preference list in $I$ of each woman in $W^{t}$ is a tie of length 2, and $M$ is complete, $w\in W^{s}$. Since $(m,w)\notin M$, $(c_{m}^{e,k},c_{w}^{e,k})$ is unassigned in $M^{\prime}$ for each $k$ ($1\leq k\leq B)$. As $(m,w)$ blocks $M$ in $I$, $w\succ_{m}M(m)$ and $m\succ_{w}M(w)$. Hence by construction, in $J$, $c_{w}^{e,k}\succ_{h_{m}}M(h_{m})$ and $c_{m}^{e,k}\succ_{h_{w}}M(h_{w})$. It follows that $|bp(J,M^{\prime})|\geq B$.

### v3

@@ 4.2 Dual markets
Suppose not. Then, $y_{4i+r}$ is either unmatched in $M^{\prime}$ or $M^{\prime}(y_{4i+r})=u^{5}_{4i+r}$. If $y_{4i+r}$ is unmatched then $(u^{5}_{4i+r},y_{4i+r})$ blocks $M^{\prime}$ in $I^{\prime}$, a contradiction. Hence, $M^{\prime}(y_{4i+r})=u^{5}_{4i+r}$ and thus $(u^{5}_{4i+r},h^{1}_{4i+r})\notin M$.
@@
We now show through the following two lemmata that if $M^{\prime}$ is a stable matching in $I^{\prime}$, then $M=M^{\prime}\setminus(U\times H)$ is a complete stable matching in $I$.
@@
Assume $z^{1}_{j}$ is unmatched in $M^{\prime}$ for some $j$ $(1\leq j\leq m)$. Thus $(p_{j}^{s},z^{2}_{j})\notin M$ $(4\leq s\leq 6)$ as $z^{2}_{j}$ must also be unmatched. Hence, $(z^{1}_{j},z^{2}_{j})$ are jointly unmatched and find $(p^{1}_{j},p^{4}_{j})$ acceptable. Furthermore, $(p^{1}_{j},p^{4}_{j})$ jointly prefers $(z^{1}_{j},z^{2}_{j})$ to any other pair. Hence, $(z^{1}_{j},z^{2}_{j})$ blocks $M$ with $(p^{1}_{j},p^{4}_{j})$, a contradiction. Thus, $z^{1}_{j}$ must be matched in any stable matching admitted by $I^{\prime}$. By a similar argument, the same holds for $z^{2}_{j}$.
@@
Let $B$ be an instance of (2,2)-e3-sat. Construct an instance $I$ of hrc-dual market as described in the proof of Theorem 22 and extend this instance as described above by adding the enforcer gadgets corresponding to the doctors in $U$ and the hospitals in $H$, to obtain the instance $I^{\prime}$ of hrc-dual market.
@@
As shown in the proof of Theorem 22, no agent outside of $Y\cup U\cup H$ can block $M^{\prime}$ in $I$. For each $i$ ($10\leq i\leq n-$ 1) and $r$ ($0\leq r\leq 3$), since $h_{4i+r}^{1}$ has its first-choice doctor in $M^{\prime}$, the couple $(u^{1}_{4i+r},u^{2}_{4i+r})$ cannot block $M^{\prime}$ in $I^{\prime}$. For a similar reason, the couple $(u^{3}_{4i+r},u^{4}_{4i+r})$ cannot block $M^{\prime}$ in $I$. Finally, $M^{\prime}(y_{4i+r})\in X$, hence $y_{4i+r}$ cannot block $M^{\prime}$ in $I^{\prime}$. Thus, $M^{\prime}$ is stable in $I^{\prime}$.
@@
Finally, the master lists shown in Figures 10, 11 and 12 indicate that the preference list of each single doctor, couple and hospital may be derived from a master list of hospital pairs, doctors and hospitals, respectively (for brevity, in the master list in each of these figures, the symbol $\succ$ is omitted from between consecutive entries). ∎
$\begin{array}[]{c}L^{1}_{i}:(y_{4i},l_{4i})~{}(c(x_{4i}),l_{4i+1})~{}(y_{4i+1},l_{4i+1})~{}(c(x_{4i+1}),l_{4i+2})~{}(c(x_{4i+3}),l_{4i+3})~{}(y_{4i+3},l_{4i+3})~{}(c(x_{4i+2}),l_{4i+2})~{}(y_{4i+2},l_{4i+2})~{}~{}~{}~{}(0\leq i\leq n-1)\\ \\ L^{2}_{i,r}:(h^{1}_{4i+r},h^{2}_{4i+r})~{}(h^{1}_{4i+r},h^{4}_{4i+r})~{}(h^{3}_{4i+r},h^{2}_{4i+r})~{}~{}~{}~{}(0\leq i\leq n-1,0\leq r\leq 3)\par~{}~{}~{}~{}~{}~{}~{}~{}~{}\par L^{3}_{j}:(z^{1}_{j},z^{2}_{j})~{}(c^{1}_{j},z^{3}_{j})~{}(c^{2}_{j},z^{4}_{j})~{}(c^{3}_{j},z^{5}_{j})~{}~{}~{}~{}(1\leq j\leq m)\\ \\ \mbox{Master List}:L^{1}_{0}~{}L^{1}_{1}~{}\dots~{}L^{1}_{n-1}~{}L^{2}_{0,0}~{}L^{2}_{0,1}~{}L^{2}_{0,2}~{}L^{2}_{0,3}~{}L^{2}_{1,0}~{}L^{2}_{1,1}~{}L^{2}_{1,2}~{}L^{2}_{1,3}~{}\ldots~{}L^{2}_{(n-1),0}~{}L^{2}_{(n-1),1}~{}L^{2}_{(n-1),2}~{}L^{2}_{(n-1),3}~{}L^{3}_{1}~{}L^{3}_{2}~{}\dots~{}L^{3}_{m}\par\end{array}$
@@
$\begin{array}[]{c}L^{4}_{j}:c^{1}_{j}~{}c^{2}_{j}~{}c^{3}_{j}~{}z^{3}_{j}~{}z^{4}_{j}~{}z^{5}_{j}\par~{}~{}~{}~{}(1\leq j\leq m)\par~{}~{}~{}~{}~{}~{}~{}~{}\par L^{5}_{i,r}:\par y_{4i+r}~{}h^{1}_{4i+r}~{}~{}~{}~{}(0\leq i\leq n-1,0\leq r\leq 3)\\ \\ \mbox{Master List}:L^{4}_{1}~{}L^{4}_{2}~{}\dots~{}L^{4}_{m}~{}L^{5}_{0,0}~{}L^{5}_{0,1}~{}L^{5}_{0,2}~{}L^{5}_{0,3}~{}L^{5}_{1,0}~{}L^{5}_{1,1}~{}L^{5}_{1,2}~{}L^{5}_{1,3}~{}\ldots~{}L^{5}_{(n-1),0}~{}L^{5}_{(n-1),1}~{}L^{5}_{(n-1),2}~{}L^{5}_{(n-1),3}\end{array}$
@@
$\begin{array}[]{c}L^{6}_{i}:x_{4i+1}~{}x_{4i}~{}x_{4i+2}~{}x_{4i+3}~{}k_{4i+3}~{}k_{4i}~{}k_{4i+2}~{}k_{4i+1}~{}~{}~{}~{}(0\leq i\leq n-1)\\ \\ L^{7}_{j}:p^{1}_{j}~{}p^{2}_{j}~{}p^{3}_{j}~{}p^{6}_{j}~{}p^{5}_{j}~{}p^{4}_{j}~{}~{}~{}~{}(1\leq j\leq m)~{}~{}~{}~{}~{}~{}~{}~{}L^{8}:s_{1}~{}s_{2}~{}\ldots~{}s_{m}~{}t_{1}~{}t_{2}~{}\ldots~{}t_{m}\\ \\ L^{9}_{i,r}:u^{5}_{4i+r}~{}u^{1}_{4i+r}~{}u^{3}_{4i+r}~{}u^{4}_{4i+r}~{}u^{2}_{4i+r}~{}~{}~{}~{}(0\leq i\leq n-1,0\leq r\leq 3)\\ \\ \mbox{Master List}:L^{7}_{1}~{}L^{7}_{2}~{}\dots~{}L^{7}_{m}~{}L^{6}_{0}~{}L^{6}_{1}~{}\dots~{}L^{6}_{n-1}~{}L^{8}~{}L^{9}_{0,0}~{}L^{9}_{0,1}~{}L^{9}_{0,2}~{}L^{9}_{0,3}~{}L^{9}_{1,0}~{}L^{9}_{1,1}~{}L^{9}_{1,2}~{}L^{9}_{1,3}~{}\ldots~{}L^{9}_{(n-1),0}~{}L^{9}_{(n-1),1}~{}L^{9}_{(n-1),2}~{}L^{9}_{(n-1),3}\end{array}$
@@ 4.3 Matchings with the minimum number of blocking pairs
min bp hrc is not approximable within $m^{1-\varepsilon}$, for any $\varepsilon>0$, where $m$ is the total length of the hospitals’ preference lists, unless P=NP. The result holds even if there are no single doctors, each couple is of type-a and finds acceptable only one pair of hospitals, and each hospital has capacity 1.
@@
Let $\varepsilon>0$ be given. We reduce from com smti, which is shown to be NP-complete in the Appendix of [26] even if every man’s list is strictly ordered and of length at most 3, and every woman’s list is either strictly ordered and of length at most 3, or is a tie of length 2. Hence, let $I$ be an instance of this restriction of com smti.
Let $U$ be the set of men in $I$ and let $W$ be the set of women in $I$. Let $G=(U\cup W,E)$ denote the underlying bipartite graph of $I$, and let $m_{I}$ denote $|E|$. Let $W^{s}$ denote the set of women in $W$, whose preference lists are strictly ordered in $I$, and let $W^{t}=W\setminus W^{s}$. Then, every woman in $W^{t}$ has a preference list in $I$ that is a tie of length 2. Let $n_{t}=|W^{t}|$, $C=\lceil\frac{3}{\varepsilon}\rceil$ and $B=n_{t}m_{I}^{C}+1$.
We construct an instance $J$ of hrc on the basis of $I$ as follows. Let $H$ be the set of hospitals in $J$, where $H=H_{U}\cup H_{W}$, $H_{U}=\{h_{u}:u\in U\}$ and $H_{W}=\{h_{w}:w\in W\}$. Let $H_{W}^{s}=\{h_{w}:w\in W^{s}\}$, and let $H_{W}^{t}=H_{W}\setminus H_{W}^{s}$.
For each person $p\in U\cup W$, let $\succeq_{p}$ denote $p$’s preference list in $I$. Each hospital $h_{p}\in H$ is initially given the preference list $\succeq_{p}$ in $J$. Note that if $p\in W^{t}$, then the tie of length 2 is broken arbitrarily in $\succeq_{p}$, so that we obtain a strictly ordered ranking $\succ_{p}$ for every hospital $h_{p}$ in $J$.
Now let $e=(u,w)\in E$. Suppose firstly that $w\in W^{t}$. Create the couple $C^{e}=(c_{u}^{e},c_{w}^{e})$ who find acceptable the hospital pair $(h_{u},h_{w})$. Replace $u$ by $c_{u}^{e}$ in $\succ_{h_{w}}$, and replace $w$ by $c_{w}^{e}$ in $\succ_{h_{u}}$. Now suppose that $w\in W^{s}$. Create the couple $C^{e,k}=(c_{u}^{e,k},c_{w}^{e,k})$ who find acceptable the hospital pair $(h_{u},h_{w})$, for each $k$ ($1\leq k\leq B$). Replace $u$ by $c_{u}^{e,1}\succ_{h_{w}}\dots\succ_{h_{w}}c_{u}^{e,B}$ in $\succ_{h_{w}}$, and replace $w$ by $c_{w}^{e,1}\succ_{h_{u}}\dots\succ_{h_{u}}c_{w}^{e,B}$ in $\succ_{h_{u}}$. Let $P(h_{w})$ denote $\succ_{h_{w}}$ at this point.
@@
Every hospital in $J$ has capacity 1. It is easy to see that there are no single doctors in $J$, and each couple is of type-a and finds acceptable only one pair of hospitals. We claim that (i) if $I$ admits a complete stable matching then $J$ admits a matching with at most $n_{t}$ blocking pairs, and (ii) if $I$ admits no complete stable matching then any matching in $J$ has at least $B$ blocking pairs.
@@
To show (i), suppose that $M$ is a complete stable matching in $I$. Construct a matching $M^{\prime}$ in $J$ as follows. For each pair $(u,w)$ in $M^{\prime}$, representing the underlying edge $e$ in $G$, suppose firstly that $w\in W^{t}$. Match the couple $(c_{u}^{e},c_{w}^{e})$ with the hospital pair $(h_{u},h_{w})$ in $M^{\prime}$. Now suppose that $w\in W^{s}$. Match the couple $(c_{u}^{e,1},c_{w}^{e,1})$ with the hospital pair $(h_{u},h_{w})$ in $M^{\prime}$. Finally, for each $h_{w}\in H_{W}$ and $k$ ($1\leq k\leq B$), match the couple $(c_{w,3}^{k},c_{w,4}^{k})$ with the hospital pair $(h_{w}^{\prime},h_{w}^{\prime\prime})$. Each hospital is then full in $M^{\prime}$, since $M$ is a complete stable matching in $I$.
@@
For each $h_{w}\in H_{W}$, and $k$ ($1\leq k\leq B$), it may be verified that $(c_{w,1}^{k},c_{w,2}^{k})$ cannot be involved in blocking $M^{\prime}$ in $J$ , since $h_{w}$ is assigned to a better doctor than $c_{i,1}^{1}$ in $M^{\prime}$. Similarly $(c_{w,5}^{k},c_{w,6}^{k})$ cannot be involved in blocking $M^{\prime}$ in $J$ for the same reason. Since $M$ is a complete stable matching in $I$, the only possible blocking pairs of $M^{\prime}$ in $J$ involve couples of the form $(c_{u}^{e},c_{w}^{e})$, where $w\in W^{t}$. Since each woman in $W^{t}$ has a preference list of length 2 in $J$, there can be at most one blocking pair for each $w\in W^{t}$, and hence $|bp(J,M^{\prime})|\leq n_{t}$.
@@
To show (ii), suppose that $I$ admits no complete stable matching, and let $M^{\prime}$ be any matching in $J$. Suppose firstly that some hospital $h_{w}$ is unassigned in $M^{\prime}$ ($h_{w}\in H_{W}$). Then $(c_{w,5}^{k},c_{w,6}^{k})$ is unassigned in $M^{\prime}$ for all $k$ ($1\leq k\leq B$). Hence, $(c_{w,5}^{k},c_{w,6}^{k})$ blocks $M^{\prime}$ with $(h_{w}^{\prime\prime},h_{w})$ for all $k$ ($1\leq k\leq B$), so $|bp(J,M^{\prime})|\geq B$. Now suppose that $(c_{w,6}^{k},h_{w})\in M^{\prime}$ for some $h_{w}\in H_{W}$ and $k$ ($1\leq k\leq B)$. Then, $(c_{w,5}^{k},h_{w}^{\prime\prime})\in M^{\prime}$. Hence, $(c_{w,1}^{k^{\prime}},c_{w,2}^{k^{\prime}})$ and $(c_{w,3}^{k^{\prime}},c_{w,4}^{k^{\prime}})$ are unassigned in $M^{\prime}$ for all $k^{\prime}$ ($1\leq k^{\prime}\leq B)$ and also $h_{w}^{\prime}$ is unassigned in $M^{\prime}$. Therefore, $(c_{w,1}^{k^{\prime}},c_{w,2}^{k^{\prime}})$ blocks $M^{\prime}$ with $(h_{w},h_{w}^{\prime})$ for all $k^{\prime}$ ($1\leq k^{\prime}\leq B$), so $|bp(J,M^{\prime})|\geq B$. Finally, suppose that $(c_{w,1}^{k},h_{w})\in M^{\prime}$ for some $h_{w}\in H_{W}$ and $k$ ($1\leq k\leq B)$. Then, $(c_{w,2}^{k},h_{w}^{\prime})\in M^{\prime}$. Hence, $(c_{w,3}^{k^{\prime}},c_{w,4}^{k^{\prime}})$ and $(c_{w,5}^{k^{\prime}},c_{w,6}^{k^{\prime}})$ are unassigned in $M^{\prime}$ for all $k^{\prime}$ ($1\leq k^{\prime}\leq B)$ and also $h_{w}^{\prime\prime}$ is unassigned in $M^{\prime}$. This implies that $(c_{w,3}^{k^{\prime}},c_{w,4}^{k^{\prime}})$ blocks $M^{\prime}$ with $(h_{w}^{\prime},h_{w}^{\prime\prime})$ for all $k^{\prime}$ ($1\leq k^{\prime}\leq B$), so $|bp(J,M^{\prime})|\geq B$.
@@
We thus assume that each $h_{w}\in H_{W}$ has an assignee better than $c_{w,1}^{1}$ in $M^{\prime}$. Let $M$ be the matching in $I$ obtained by replacing each assignment of a couple $(c_{u}^{e},c_{w}^{e})$ with the pair $(u,w)$ in $M$, and by replacing each assignment of a couple $(c_{u}^{e,k},c_{w}^{e,k})$ with the pair $(u,w)$ in $M$. It is easy to see that $M$ is a matching in $I$, since each of $(c_{u}^{e},c_{w}^{e})$ and $(c_{u}^{e,k},c_{w}^{e,k})$ finds acceptable only $(h_{u},h_{w})$, and each hospital has capacity 1 in $J$. Moreover, $M$ is a complete matching in $I$, since in $M^{\prime}$, each hospital $h_{w}\in H_{W}$ has an assignee better than $c_{w,1}^{1}$. Hence by assumption, $M$ is not stable in $I$, so $M$ admits a blocking pair $(u,w)$ in $I$.
@@
Let $e=(u,w)\in E$. Since the preference list in $I$ of each woman in $W^{t}$ is a tie of length 2, and $M$ is complete, $w\in W^{s}$. Since $(u,w)\notin M$, $(c_{u}^{e,k},c_{w}^{e,k})$ is unassigned in $M^{\prime}$ for each $k$ ($1\leq k\leq B)$. As $(u,w)$ blocks $M$ in $I$, $w\succ_{u}M(u)$ and $u\succ_{w}M(w)$. Hence by construction, in $J$, $c_{w}^{e,k}\succ_{h_{u}}M(h_{u})$ and $c_{u}^{e,k}\succ_{h_{w}}M(h_{w})$. It follows that $|bp(J,M^{\prime})|\geq B$.